OpenStudy (anonymous):

Rueben's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Rueben $4.40 per pound, and type B coffee costs $5.70 per pound. This month, Rueben made 165 pounds of the blend, for a total cost of $841.70 . How many pounds of type B coffee did he use?

OpenStudy (anonymous):

165=x+y 841.70=4.4x+5.7y

OpenStudy (anonymous):

you can see this by making pounds of type A = x and pounds of type B = y you know he bought 165 pounds therefore you know that x and y have to equal 165 165=x+y now you know it costed 841.70 all together and you know that type a costs 4.4 per pound(x) and type b costs 5.7 per pound (y)... replacing these into a equation you get 841.7=4.4x+5.7y

OpenStudy (anonymous):

so how do you get the answer

OpenStudy (anonymous):

now use substitution by solving the top equation for y 165=x+y 165-x=x-x+y [subtract both sides by x] 165-x=y take this equation and substitute it into the second 841.70=4.4x+5.7(165-x) 841.70=4.4x+940.5-5.7x [distributed 5.7 to (165-x)] 841.70-940.5=4.4x+940.5-940.5-5.7x [ subtracted 940.5 from both sides] -98.8=-1.3x [added like terms] -98.8/-1.3=-1.3x/-1.3 [divided by -1.3 on both sides] 76 = x now you know that he bought 76 pounds of Type A to find Type B go back to your first equation 165=x+y since x=76 165=76+y 165-76=76-76+y [substract 76 from both sides] 89= y 89 bounds of type B

OpenStudy (mertsj):

Let x =amount of Type A then 165-x= amount of Type B 4.40x +5.70(165-x)=841.70

OpenStudy (mertsj):

Solve that equation.